OpenStudy (leozap1):

Fan & Medal... Walker is buying a house for $225,000. He is financing $195,000 and obtained a 30-year, fixed-rate mortgage with a 6.625% interest rate. How much are his monthly payments? $1,248.61 $1,076.56 $1,712.09 $1,440.70

OpenStudy (photon336):

the interest rate is yearly?

OpenStudy (leozap1):

It doesn't say, but if he pays monthly, I would guess that the interest would be monthly.

OpenStudy (leozap1):

Wouln't the interest be on the $195,000 since he is financing it?

OpenStudy (photon336):

yeah, I think you're right. What I think we can do is probably use the simple interest formula \[\frac{ P*R*T }{ 100 } = Simple~interest \]

OpenStudy (leozap1):

When I used the amount $195,000 I ended up getting $ 1,242.19.

OpenStudy (photon336):

what formula did you use

OpenStudy (leozap1):

I used this, I think it's the same formula you said. http://www.calculatorsoup.com/calculators/financial/interest-only-mortgage-calculator.php

OpenStudy (leozap1):

Nvm, I see what happened when I just used the formula you posted instead of the calculator. I got 1,440.70 .

OpenStudy (photon336):

remember that the interest rate is usually anaully

OpenStudy (leozap1):

Yea that's how I did it, because monthly gave me a wrong answer.

OpenStudy (photon336):

yeah so put in 30 years first.

OpenStudy (photon336):

Yeah for some reason i'm getting this. \[\frac{ 195,000*6.625*30 }{ 100 } = 387,562.5/(360) = 1076.56 \]

OpenStudy (leozap1):

Idk, I submitted the quick check and it turned out to be $1,248.61.